# Facilitating Long Context Understanding via Supervised Chain-of-Thought Reasoning

### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have enabled them to process increasingly longer sequences, ranging from 2K to 2M tokens and even beyond. However, simply extending the input sequence length does not necessarily lead to effective long-context understanding. In this study, we integrate Chain-of-Thought (CoT) reasoning into LLMs in a supervised manner to facilitate effective long-context understanding. To achieve this, we introduce LongFinanceQA, a synthetic dataset in the financial domain designed to improve long-context reasoning. Unlike existing long-context synthetic data, LongFinanceQA includes intermediate CoT reasoning before the final conclusion, which encourages LLMs to perform explicit reasoning, improving accuracy and interpretability in long-context understanding. To generate synthetic CoT reasoning, we propose Property-driven Agentic Inference (PAI), an agentic framework that simulates human-like reasoning steps, including property extraction, retrieval, and summarization. We evaluate PAI’s reasoning capabilities by assessing GPT-4o-mini w/ PAI on the Loong benchmark, outperforming standard GPT-4o-mini by 20.0%. Furthermore, we fine-tune LLaMA-3.1-8B-Instruct on LongFinanceQA, achieving a 24.6% gain on Loong’s financial subset.
@@
Jingyang Lin1,2 Andy Wong2, Tian Xia2, Shenghua He2, Hui Wei2,3, Mei Han2, Jiebo Luo1 1 University of Rochester, NY, United States 2 PAII Inc., CA, United States 3 University of California, Merced, CA, United States
footnotetext: Corresponding author: ✉ jlin81@ur.rochester.edu.footnotetext: Work was done during the internship at PAII Inc.
@@ 1 Introduction
To validate this hypothesis, we introduce LongFinanceQA, a novel long-context synthetic dataset constructed using financial data. Each sample in this dataset consists of a practical long-context question and the corresponding augmented answer along with intermediate chain-of-thought (CoT) reasoning steps. In particular, we first collect 6,911 bilingual financial annual reports (i.e., English and Chinese) published before 2022. We then build a financial metric pool comprising key metrics commonly found in these reports, such as profit, cash flow, and debt. Based on these financial metrics, we generate 46,457 long-context questions that require single- or multi-source evidence. To generate reliable reasoning-augmented answers, we propose Property-driven Agentic Inference (PAI), a comprehensive agentic framework. PAI leverages LLM-based agents to simulate human-like reasoning and operates in three steps: 1) a property extraction agent decomposes complex queries by identifying key properties, where each property consists of a metric and the corresponding subject; 2) a property-based retrieval agent retrieves relevant information from long documents for each identified property, and then generates intermediate findings by leveraging the property-based retrieved content; 3) a summarization agent synthesizes the intermediate findings into a coherent conclusion. The synthetic reasoning results are formed by integrating outputs from the property extraction and retrieval stages, while the conclusion is derived from the summarization stage. Finally, LongFinanceQA comprises 46,457 long-context QA pairs with CoT reasoning over 6,911 financial reports.
@@
Empirically, we first evaluate the outcome of PAI on the Loong benchmark Wang et al. (2024), involving challenging long-context tasks on three different domains. The results show that equipping GPT-4o-mini with PAI achieves a substantial 20% improvement on the Loong as shown in Figure 1. Moreover, the effectiveness of PAI guarantees the quality of the synthetic data in LongFinanceQA. Meanwhile, the enhanced LLaMA-3.1 model, LongPAI, achieves a 24.6% improvement on the Financial subset of Loong. Notably, in several scenarios, LongPAI even surpasses its teacher model PAI. This phenomenon emphasizes the importance of long-context modeling, contradicting recent arguments that the long-context problem can be solved by short language models Qian et al. (2024); Chen et al. (2024b).
Our main contributions are three-fold: 1) we introduce LongFinanceQA, a long-context synthetic dataset for fine-tuning with 46,457 QA pairs featuring high-quality CoT reasoning from 6,911 bilingual financial annual reports; 2) to generate reasoning-augmented answers, we propose an agentic framework, Property-driven Agentic Inference, to mimic human behaviors; and 3) empirical results validate the effectiveness of PAI and supervised CoT reasoning, the quality of LongFinanceQA, and the importance of long-context modeling.
@@ 3.1 Problem Formulation
where $\theta$ is the parameters of language models.
@@ 3.2 LongFinanceQA Dataset
Augmented Answer Generation with CoT Reasoning. Given long-context questions and their corresponding documents, our goal is to generate answers with CoT reasoning, following the principle of Eq. (3). In particular, we first generate step-by-step reasoning based on the question and documents, then integrate the reasoning steps into a conclusion. Generally, long-context questions are challenging as they require models to retrieve multiple pieces of evidence scattered throughout long content and then integrate them for a global understanding Wang et al. (2024); Edge et al. (2024). Inspired by this phenomenon, we intend to first extract key evidence points that support answering the question, then retrieve relevant information based on these points, and finally aggregate them into the conclusion. We term these supporting evidence points as “Properties”. Following this methodology, we propose the Property-driven Agentic Inference (PAI) framework, illustrated in Figure 2, containing three steps: (1) property extraction, (2) property-based retrieval, and (3) summarization.
[…]
@@ 3.3 Supervised CoT Reasoning
where $N_{r}$ is the number of reasoning tokens and $N_{y}$ indicates the answering token, including those in the properties, sub-answers, and final answers. We mask non-answer positions during the fine-tuning, allowing for more efficient learning. Although we only consider the LLaMA-3.1 model here, LongFinanceQA can also be used to fine-tune other language models. Please refer to Section 4.1 for more implementation details.
@@ 4 Experiments
LongPAI (ours) 262K 77.89 0.63 62.00 0.45 48.00 0.17 51.33 0.47 60.36 0.41
4.1 Experimental Setup
[…]
Base Models. We adopt GPT-4o-mini Achiam et al. (2023) and LLaMA-3.1-8B-Instruct Dubey et al. (2024) as our base models. Specifically, GPT-4o-mini serves as the agent of the proposed Preference Agentic Inference (PAI), while LLaMA-3.1-8B is used as the base model for LongPAI, which is fine-tuned on the LongFinanceQA dataset.
Implementation Details. To enable training on long sequences (> 250K), we employ several optimization techniques, including flash-attention-2 Dao (2023) and ring-attention Liu et al. (2024c). Furthermore, we adopt a zigzag sharding approach Zhu (2024) within ring attention for more effective load distribution across multiple GPUs. This training setup allows us to fine-tune the large language models fully. Using 8 A100 GPUs, the long-context training is completed in three days for fine-tuning. In addition, in training the long-context LLMs, we adopt the rotary base scaling approach Liu et al. (2024d) and scale up the base value (i.e., rotary base of 1,247,820) to adapt RoPE to a longer context. For optimization, we use a constant learning rate of 2e-6 for the entire training procedure. Following the common practice Fu et al. (2024), we set the batch size to 16M tokens as mentioned in Dubey et al. (2024). During inference, we set the temperature as zero to eliminate the randomness. We also increase the maximum output tokens to 1,024 since the CoT reasoning requires more output tokens.
4.2 Main Results
[…]
@@ 4.2 Main Results
LongPAI (ours) 32.46 32.38
Table 4: Ablation study on Supervised CoT Reasoning. Symbol § represents LongPAI without supervised CoT reasoning during fine-tuning. Average Scores (0-100) are evaluated by GPT-4-Turbo. Green highlights the remarkable improvements over the base GPT-4o-mini, while Red indicates a decline. Abbreviations: S.L. (Spotlight Locating), Comp. (Comparison), Clust. (Clustering), and Chain. (Chain of Reasoning).
@@
LongPAI 77.89 62.00 48.00 51.33 60.36
4.3 Discussion
Ablation Study on Supervised CoT Reasoning. To further analyze the impact of supervised CoT reasoning, we fine-tune the base LLaMA-3.1 model on LongFinanceQA while excluding CoT reasoning steps from the augmented answers, resulting in a new model, LongPAI§. Unlike the original LongPAI, LongPAI§ directly predicts the final answer, skipping intermediate reasoning steps. Table 4 shows that LongPAI significantly outperforms LongPAI§ over different input lengths. This result strongly supports the hypothesis mentioned in Section 1, which argues that directly guiding models to generate brief answers without intermediate reasoning steps for long-context modeling will lead to suboptimal training (finding 1). Moreover, Table 4 presents several interesting finding. First, while LongPAI§ performs comparably to LongPAI on short content (10K–50K tokens), its performance declines significantly on longer content, which means CoT reasoning is necessary for long-context modeling (finding 2). Furthermore, LongPAI§ performs well in single-source questions (Spotlight Locating) but struggles with multi-source questions (Comparison, Clustering, and Chain of Reasoning) as input length increases, demonstrating that CoT reasoning benefits complex long-context problem-solving (finding 3). In sum, all these findings reaffirm the importance of the reasoning capability for long-context modeling.
Comparison of Various Inference Frameworks. We compare PAI with two relevant inference frameworks: PAI- and RAG. PAI- is a variant of PAI that generates sub-questions directly, rather than first extracting properties and then generating sub-questions. Table 5 presents that PAI- generally outperforms the base GPT-4o-mini, except on $\mathtt{Set1}$. However, there is a gap between PAI- and PAI, highlighting the superiority of the Property Extraction Agent. Retrieval-Augmented Generation (RAG) Lewis et al. (2020) follows a two-step process: it first retrieves the top $K$ chunks relevant to a given query, and then uses retrieved chunks to generate an answer. Following Loong Wang et al. (2024), we adopt the BGE Chen et al. (2024a) as the embedding choice and set $K$ to 50, selecting from 5, 10, 30, and 50. Table 5 shows that RAG struggles with long-context problems, a conclusion also reached by Wang et al. (2024).
@@ 4.3 Discussion
Efficiency Analysis on PAI. The PAI framework relies on multi-step inference, whereas LongPAI achieves results with a single inference step. We adopt the number of input tokens processed on the Loong benchmark (i.e., 1,600 samples) as a metric to compare the efficiency between PAI and LongPAI. In this comparison, PAI processes 3.53B tokens in total, determined by the GPT-4o tokenizer, whereas LongPAI requires only 112M tokens—less than 3% of PAI’s total. Despite PAI delivering the strongest overall performance, LongPAI stands out as the far more efficient approach, demonstrating a dramatic reduction in computational cost.
5 Conclusion
In this work, we introduce LongFinanceQA, a novel long-context synthetic dataset featuring reasoning-augmented answers for practical long-context questions. To generate these enriched answers, we develop the Property-driven Agentic Inference (PAI) framework, which simulates human-like reasoning through property extraction, retrieval, and summarization. Empirical analysis confirms the effectiveness of PAI and ensures the high quality of LongFinanceQA. Beyond PAI, we explicitly guide a language model in learning CoT reasoning using LongFinanceQA. A series of empirical results underscore the importance of reasoning-augmented long-context modeling.
@@ 5 Conclusion
In this study, we investigated the effectiveness of supervised CoT reasoning using reasoning-augmented long-context synthetic data (LongFinanceQA). While effective, our approach has certain limitations. First, although the reasoning-enhanced LongPAI model demonstrates significant improvement in the financial domain, its ability to generalize to broader long-context scenarios remains uncertain. To address this, our future work will explore the impact of diverse data sources and different data scales. Second, while the PAI framework achieves strong performance on the Loong benchmark, it still relies on human-crafted prompts to guide agents in maintaining structured reasoning. Therefore, we plan to explore inference methods that enable autonomous reasoning with long-context inputs, minimizing reliance on extensive human-crafted prompts.

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have enabled them to process increasingly longer sequences, ranging from 2K to 2M tokens and even beyond. However, simply extending the input sequence length does not necessarily lead to effective long-context understanding. In this study, we integrate Chain-of-Thought (CoT) reasoning into LLMs in a supervised manner to facilitate effective long-context understanding. To achieve this, we introduce LongFinanceQA, a synthetic dataset in the financial domain designed to improve long-context reasoning. Unlike existing long-context synthetic data, LongFinanceQA includes intermediate CoT reasoning before the final conclusion, which encourages LLMs to perform explicit reasoning, improving accuracy and interpretability in long-context understanding. To generate synthetic CoT reasoning, we propose Property-based Agentic Inference (PAI), an agentic framework that simulates human-like reasoning steps, including property extraction, retrieval, and summarization. We evaluate PAI’s reasoning capabilities by assessing GPT-4o-mini w/ PAI on the Loong benchmark, outperforming standard GPT-4o-mini by 20.0%. Furthermore, we fine-tune LLaMA-3.1-8B-Instruct on LongFinanceQA, achieving a 28.0% gain on Loong’s financial subset.
@@
Jingyang Lin1,2 Andy Wong2, Tian Xia2, Shenghua He2, Hui Wei2,3, Mei Han2, Jiebo Luo1 1 University of Rochester, NY, United States 2 PAII Inc., CA, United States 3 University of California, Merced, CA, United States https://long-pai.github.io/
footnotetext: Corresponding author: 🖂 jlin81@ur.rochester.edu.footnotetext: Work was done during the internship at PAII Inc.
@@ 1 Introduction
To validate this hypothesis, we introduce LongFinanceQA, a novel long-context synthetic dataset constructed using financial data. Each sample in this dataset consists of a practical long-context question and the corresponding augmented answer along with intermediate chain-of-thought (CoT) reasoning steps. In particular, we first collect 6,911 bilingual financial annual reports (i.e., English and Chinese) published before 2022. We then build a financial metric pool comprising key metrics commonly found in these reports, such as profit, cash flow, and debt. Based on these financial metrics, we generate 46,457 long-context questions that require single- or multi-source evidence. To generate reliable reasoning-augmented answers, we propose Property-based Agentic Inference (PAI), a comprehensive agentic framework. PAI leverages LLM-based agents to simulate human-like reasoning and operates in three steps: 1) a property extraction agent decomposes complex queries by identifying key properties, where each property consists of a metric and the corresponding subject; 2) a property-based retrieval agent retrieves relevant information from long documents for each identified property, and then generates intermediate findings by leveraging the property-based retrieved content; 3) a summarization agent synthesizes the intermediate findings into a coherent conclusion. The synthetic reasoning results are formed by integrating outputs from the property extraction and retrieval stages, while the conclusion is derived from the summarization stage. Finally, LongFinanceQA comprises 46,457 long-context QA pairs with CoT reasoning over 6,911 financial reports.
@@
Empirically, we first evaluate the outcome of PAI on the Loong benchmark Wang et al. (2024), involving challenging long-context tasks on three different domains. The results show that equipping GPT-4o-mini with PAI achieves a substantial 20% improvement on the Loong as shown in Figure 1. Moreover, the effectiveness of PAI guarantees the quality of the synthetic data in LongFinanceQA. Meanwhile, the enhanced LLaMA-3.1 model, LongPAI, achieves a 28.0% improvement on the Financial subset of Loong. Notably, in several scenarios, LongPAI even surpasses its teacher model PAI. This phenomenon emphasizes the importance of long-context modeling, contradicting recent arguments that the long-context problem can be solved by short language models Qian et al. (2024); Chen et al. (2024b).
Our main contributions are three-fold: 1) we introduce LongFinanceQA, a long-context synthetic dataset for fine-tuning with 46,457 QA pairs featuring high-quality CoT reasoning from 6,911 bilingual financial annual reports; 2) to generate reasoning-augmented answers, we propose an agentic framework, Property-based Agentic Inference, to mimic human behaviors; and 3) empirical results validate the effectiveness of PAI and supervised CoT reasoning, the quality of LongFinanceQA, and the importance of long-context modeling.
@@ 2 Related Work
Agentic RAG. Agentic RAG frameworks can be grouped into two paradigms Liang et al. (2025): predefined reasoning and agentic reasoning RAG. Our proposed PAI framework is a hybrid of the two paradigms. Like predefined RAG systems Press et al. (2023); Sarthi et al. (2024), PAI follows a structured, multi-step reasoning workflow. However, the components of PAI employ the techniques used in agentic reasoning RAG. Concretely, Property Extraction Agent and Proper-based Retrieval Agent use prompt-based approaches (i.e., function calling Eleti et al. (2023)) to generate structured outputs. Agentic RAG systems Li et al. (2024); Yao et al. (2023); Eleti et al. (2023) typically treat follow-up queries as free-form text, which invites query drift, irrelevant results, and hard-to-audit reasoning. Different from agentic RAG, PAI instead converts each sub-query into a well-formed property, enabling verifiable retrieval, interpretable reasoning traces, and compact supervised training. Beyond agentic RAG, this study further investigates transferring the capability of long-context reasoning into a lightweight language model via supervised CoT, resulting in the LongPAI model.
@@ 3 Methodology
Figure 2: Overview of Property-based Agentic Inference (PAI), containing three stages. A Property Extraction Agent identifies key properties $\mbox{$p$}_{i}$ from the given query $\mathbf{Q}$, where each property consists of a measurable metric and its corresponding subject. Given the selected properties, a Property-based Retrieval Agent first transforms each property into a sub-query $\mbox{$q$}_{i}$ to retrieve relevant content chunks from long documents, yielding intermediate findings. A Summarization Agent integrates these intermediate findings to generate a comprehensive conclusion $\mathbf{A}$. After finishing PAI, we incorporate the output from the above three agents to produce reasoning-augmented answers. These augmented answers serve as the core contribution of the LongFinanceQA.
@@ 3.1 Problem Formulation
where $\theta$ is the parameter of the language models.
@@ 3.2 LongFinanceQA Dataset
Augmented Answer Generation with CoT Reasoning. Given long-context questions and their corresponding documents, our goal is to generate answers with CoT reasoning, following the principle of Eq. (3). In particular, we first generate step-by-step reasoning based on the question and documents, then integrate the reasoning steps into a conclusion. Generally, long-context questions are challenging as they require models to retrieve multiple pieces of evidence scattered throughout long content and then integrate them for a global understanding Wang et al. (2024); Edge et al. (2024). Inspired by this phenomenon, we intend to first extract key evidence points that support answering the question, then retrieve relevant information based on these points, and finally aggregate them into the conclusion. We term these supporting evidence points as “Properties”. Following this methodology, we propose the Property-based Agentic Inference (PAI) framework, illustrated in Figure 2, containing three steps: (1) property extraction, (2) property-based retrieval, and (3) summarization.
Step 1: Property Extraction. This step focuses on extracting a set of properties $\{\mbox{$p$}_{i}\}^{N_{p}}_{i=1}$ from the given query $\mathbf{Q}$, where $N_{p}$ denotes the number of properties. Each property consists of a metric (a measurable factor being analyzed) and its corresponding subject mentioned in the query $\mathbf{Q}$. For instance, given the query shown in Figure 2, the metric is “revenue” and subjects are “<Company A>’s annual reports from different years (e.g., 2018, 2020, and 2022)”, which serve as the sources of information to determine the trend.
Step 2: Property-based Retrieval. After extracting the properties $\{\mbox{$p$}_{i}\}^{N_{p}}_{i=1}$ from the given query, this step aims to retrieve relevant information based on these properties. Specifically, each property is first transformed into a sub-query $\mbox{$q$}_{i}$, which is then matched against relevant content chunks. These chunks are derived from the original long documents $\mathbf{X}$, with each chunk limited to 1,024 tokens. Based on the sub-queries and their corresponding retrieved chunks, we can derive several intermediate findings (i.e., sub-answers).
@@ 3.3 Supervised CoT Reasoning
where $N_{r}$ is the number of reasoning tokens and $N_{y}$ indicates the answering token, including those in the properties, sub-answers, and final answers. We mask non-answer positions during the fine-tuning, allowing for more efficient learning. Although we only consider the LLaMA-3.1 model here, LongFinanceQA can also be used to fine-tune other language models (See Section 4.1 for details).
@@ 4.1 Experimental Setup
[…]
Base Models. We adopt GPT-4o-mini Achiam et al. (2023) and LLaMA-3.1-8B-Instruct Dubey et al. (2024) as our base models. Specifically, GPT-4o-mini serves as the agent of the proposed Property-based Agentic Inference (PAI), while LLaMA-3.1-8B is used as the base model for LongPAI, which is fine-tuned on the LongFinanceQA.
Implementation Details. To enable training on long sequences (> 250K), we employ several optimization techniques, including flash-attention-2 Dao (2023) and ring-attention Liu et al. (2024c). Furthermore, we adopt a zigzag sharding approach Zhu (2024) within ring attention for more effective load distribution across multiple GPUs. This training setup allows us to fine-tune the large language models fully. Using 8 A100 GPUs, the long-context training is completed in three days for fine-tuning. In addition, in training the long-context LLMs, we adopt the rotary base scaling approach Liu et al. (2024d) and scale up the base value (i.e., rotary base of 1,247,820) to adapt RoPE to a longer context. For optimization, we use a constant learning rate of 1e-5 for the entire training procedure. Following the common practice Fu et al. (2024), we set the batch size to 16M tokens as mentioned in Dubey et al. (2024). During inference, we set the temperature as zero to eliminate the randomness. We also increase the maximum output tokens to 1,024 since the CoT reasoning requires more output tokens.
4.2 Main Results
Human Evaluation on LongFinanceQA. To further quantify the data quality, we manually annotated 200 randomly selected LongFinanceQA questions, including both single-source and multi-source questions. Then, 10 volunteers score each response on a 1-to-5 scale (1 = completely incorrect, 5 = completely correct). Furthermore, volunteers rate the correctness of intermediate reasoning steps on a 1-to-5 scale for each multi-source question. The averaged scores above 4.0 can be interpreted as largely correct. Table 2 shows that both final answers and their supporting reasoning trajectories are consistently reliable. It provides concrete evidence that the performance improvements primarily arise from high-quality synthesized data, rather than from patterns in noisy samples.
[…]
Table 2: Human evaluation on LongFinanceQA dataset.
Subset Score
Single-source QA 4.45 $\pm$ 0.21
Multi-source QA 4.24 $\pm$ 0.23
Intermediate Reasoning Trajectories 4.30 $\pm$ 0.22
Overall 4.29 $\pm$ 0.19
@@ 4.2 Main Results
LongPAI (ours) 262K 71.48 0.59 56.50 0.45 46.83 0.17 76.00 0.67 60.92 0.43
@@
LongPAI (ours) 32.65 32.88
4.3 Discussion
Ablation Study on Supervised CoT Reasoning. To further analyze the impact of supervised CoT reasoning, we fine-tune the base LLaMA-3.1 model on LongFinanceQA while excluding CoT reasoning steps from the augmented answers, resulting in a new model, LongPAI§. Unlike the original LongPAI, LongPAI§ directly predicts the final answer, skipping intermediate reasoning steps. Table 5 shows that LongPAI significantly outperforms LongPAI§ over different input lengths. This result strongly supports the hypothesis mentioned in Section 1, which argues that directly guiding models to generate brief answers without intermediate reasoning steps for long-context modeling will lead to suboptimal training (finding 1). Moreover, Table 5 presents several interesting findings. First, while LongPAI§ performs comparably to LongPAI on short content (10K–50K tokens), its performance declines significantly on longer content, which means CoT reasoning is necessary for long-context modeling (finding 2). Furthermore, LongPAI§ performs well in single-source questions (Spotlight Locating) but struggles with multi-source questions (Comparison, Clustering, and Chain of Reasoning) as input length increases, demonstrating that CoT reasoning benefits complex long-context problem-solving (finding 3). In sum, all these findings reaffirm the importance of the reasoning capability for long-context modeling.
Comparison of Various Inference Frameworks. We compare PAI with four relevant inference frameworks: PAI-, Self-Route Li et al. (2024), Self-Ask Press et al. (2023), and RAG Lewis et al. (2020). PAI- is a variant of PAI that generates sub-questions directly, rather than first extracting properties and then generating sub-questions. Table 6 presents that PAI- generally outperforms the base GPT-4o-mini, except on $\mathtt{Set1}$. However, there is a gap between PAI- and PAI, highlighting the superiority of the Property Extraction Agent. RAG follows a two-step process: it first retrieves the top $K$ chunks relevant to a given query, and then uses the retrieved chunks to generate an answer. Following Loong Wang et al. (2024), we adopt the BGE Chen et al. (2024a) as the embedding choice and set $K$ to 50, selecting from 5, 10, 30, and 50. Table 6 shows that RAG struggles with long-context problems, a conclusion also reached by previous work Wang et al. (2024). Furthermore, we compare the proposed PAI framework with both predefined reasoning RAG (Self-Route) and agentic reasoning RAG (Self-Ask). The results demonstrate that PAI consistently outperforms the other two agentic RAG methods, especially on longer contexts (Set3 and Set4), highlighting a significant superiority of the PAI framework under the scenario of long-context understanding.
Table 5: Ablation study on Supervised CoT Reasoning. Symbol § represents LongPAI without supervised CoT reasoning during fine-tuning. Average Scores (0-100) are evaluated by GPT-4-Turbo. Green highlights the remarkable improvements over the base LLaMA-3.1-8B, while Red indicates a decline. Abbreviations: S.L. (Spotlight Locating), Comp. (Comparison), Clust. (Clustering), and Chain. (Chain of Reasoning).
@@ 4.3 Discussion
LongPAI 71.48 56.50 46.83 76.00 60.92
Efficiency Analysis on PAI and LongPAI. The PAI framework relies on multi-step inference, whereas LongPAI achieves results with a single inference step. We adopt the number of input tokens processed on the Loong benchmark (i.e., 1,600 samples) as a metric to compare the efficiency between PAI and LongPAI. In this comparison, PAI processes 3.53B tokens in total, determined by the GPT-4o tokenizer, whereas LongPAI requires only 112M tokens, less than 3% of PAI’s total. Despite PAI delivering the highest overall performance, LongPAI stands out as the far more efficient approach, demonstrating a dramatic reduction in computational cost.
5 Conclusion
In this work, we introduce LongFinanceQA, a novel long-context synthetic dataset featuring reasoning-augmented answers for practical long-context questions. To generate these enriched answers, we develop the Property-based Agentic Inference (PAI) framework, which simulates human-like reasoning through property extraction, retrieval, and summarization. Empirical analysis demonstrates the effectiveness of PAI and ensures the high quality of the LongFinanceQA dataset. Beyond PAI, we explicitly guide a lightweight language model in learning CoT reasoning via fine-tuning on LongFinanceQA. A series of empirical results underscores the importance of reasoning-augmented long-context modeling.
@@ 5 Conclusion
In this study, we investigate the effectiveness of supervised CoT reasoning using reasoning-augmented long-context synthetic data (LongFinanceQA). While effective, our approach has certain limitations. First, although the reasoning-enhanced LongPAI model demonstrates significant improvement in the financial domain, its ability to generalize to broader long-context scenarios remains uncertain since overfitting is an inherent issue that stems from supervised fine-tuning Li et al. (2025). To address this, our future work will explore the impact of diverse data sources and different data scales. Also, we might explore potential techniques to solve this issue, such as regularization Li et al. (2025) and reinforcement learning Schulman et al. (2017); Shao et al. (2024); Rafailov et al. (2023) Second, while the PAI framework achieves strong performance on the Loong benchmark, it still relies on domain-specific human-crafted prompts to guide agents in maintaining structured reasoning. For example, prompts may explicitly direct the model to focus on concepts such as “profit, revenue, or cash flow” when processing financial data, or “reference” and “citation” when handling scientific texts. Therefore, we plan to explore inference methods that enable autonomous reasoning with long-context inputs, minimizing reliance on extensive human-crafted prompts.
Despite the current limitations of PAI and LongPAI, their ability to achieve significant performance gains through minimal domain-specific guidance highlights their strong practical potential for specialized applications.